What are the solutions to (x + 6)^2 = 36?
A. 6 and 0
B. -12 and 0
C. 6 and 36
D. 42 and 30

Found 3 solutions by josgarithmetic, ikleyn, greenestamps:
Answer by josgarithmetic(39630)

what trouble do you have with this equation?

Answer by ikleyn(52879)

.

     Step by step



    (x + 6)^2 = 36


Take square root of both sides

    x+6 = +/- 6.


Consider two cases in succession


    (a)  x+6 = 6  --->    x= 0.


    (b)  x+6 = -6  --->  x= -12.


ANSWER.  The two solutions are -12 and 0.

Solved.

---------------

This case is an example of a quadratic equation, which can be easily solved
without using the quadratic formula.



Answer by greenestamps(13209)



(x+6)^2 = 36
x^2+12x+36 = 36
x^2+12x = 0
x(x+12) = 0

x = 0

or

x+12 = 0
x = -12

ANSWERS: 0 and -12